# Simultaneous false discovery bounds for invariant causal prediction

### v2

@@ Abstract
Invariant causal prediction (ICP, Peters et al., (2016)) provides a novel way to identify causal predictors of a response by utilizing heterogeneous data from different environments. One advantage of ICP is that it guarantees to make no false causal discoveries with high probability. Such a guarantee, however, can be too conservative in some applications, resulting in few or no discoveries. To address this, we propose simultaneous false discovery bounds for ICP, which provides users with extra flexibility in exploring causal predictors and can extract more informative results. These additional inferences come for free, in the sense that they do not require additional assumptions, and the same information obtained by the original ICP is retained. We demonstrate the practical usage of our method through simulations and a real dataset.
@@ 1 Introduction
Discovering causal predictors of the response of interest is usually the primary goal of scientific research. Based solely on observational data, the causal relationship might be unidentifiable. In such cases, heterogeneous data from different environments can be helpful for further identifiability and thus is a valuable source for causal discovery. Under the multi-environments setting, Peters et al., (2016) proposed a novel method, called invariant causal prediction (ICP), to identify causal predictors of a response by exploiting the invariance of the conditional distribution of the response given its causal predictors across environments. Compared to other causal discovery algorithms, one main advantage of ICP is that it provides a statistical confidence guarantee for its output $\widehat{S}^{\text{ICP}}$:
$P\left(\widehat{S}^{\text{ICP}}\subseteq S^{*}\right)\geq 1-\alpha,$
@@
where $\alpha\in(0,1)$ is a nominal level and $S^{*}$ denotes the set of true causal predictors. That is, ICP ensures that all its discoveries are true causal predictors with a probability larger than $1-\alpha$. This is also known as the familywise error rate control guarantee.
The original ICP approach has been generalized to non-linear setting (Heinze-Deml et al.,, 2018), sequential data (Pfister et al.,, 2019), and transformation models (Kook et al.,, 2023). All these methods inherit the familywise error rate control guarantee. However, controlling the familywise error rate can be too conservative in many applications, especially when causal predictors are highly correlated with non-causal ones. In such cases, ICP may result in few or no causal discoveries. To address this, Heinze-Deml et al., (2018) proposed the so-called defining sets, which are the smallest sets guaranteeing to contain at least one true causal predictor. They are especially useful in the case where ICP returns an empty set.
@@
In this paper, we propose a method to obtain simultaneous false discovery bound (Genovese and Wasserman,, 2006; Goeman and Solari,, 2011; Goeman et al.,, 2021) for ICP. Specifically, let $\mathcal{N}^{*}\subseteq[m]=\{1,\dots,m\}$ be the index set of non-causal predictors among $m$ predictors, the simultaneous false discovery bound is a function $t_{\alpha}:2^{[m]}\rightarrow\{0,\dots,m\}$, where $2^{m}$ denotes the power set of $[m]$, such that
$P(|R\cap\mathcal{N}^{*}|\leq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$
As a result, one can freely check any set $R$ of interest, and a high probability upper bound of the false discoveries can be obtained by $t_{\alpha}(R)$. This contains the familywise error rate control and the defining sets of Heinze-Deml et al., (2018) as special cases. In particular, by searching for the largest set $R$ whose corresponding $t_{\alpha}(R)$ is 00, we obtain a set with familywise error rate control; by searching for the smallest sets whose corresponding $t_{\alpha}(R)\leq|R|-1$, we obtain all defining sets.
@@
Extra causal information can be extracted using the simultaneous false discovery bound. As a quick illustration, for a simulated dataset with nine predictors shown in Section 4.1, no information is obtained by the original ICP approach as it returns zero discoveries. However, more causal information can be extracted by using our simultaneous false discovery bounds. For example, we can know that with a probability larger than $0.95$, there is at least one causal predictor in $\{X_{3},X_{7}\}$, at least two causal predictors in $\{X_{3},X_{5},X_{7},X_{8}\}$, and at least four causal predictors in all nine predictors. More simulations and a real data application can be found in Section 4.1.
To obtain the simultaneous false discovery bound for ICP, we proceed by considering a multiple testing problem (see (6)) that directly tests causal predictors. We show that the original ICP approach is equivalent to directly comparing certain p-values (see (8)) to the significance level. Then, by generalizing the null hypothesis (6) and p-value (7) to (10) and (11), respectively, we propose a simultaneous false discovery upper bound (12). This upper bound can be seen as a special case of the closed testing method by Goeman and Solari, (2011), but in our case, no closed testing adjustment is needed due to the specific form of the p-values. In comparison to the original ICP approach, our method provides additional causal information at the cost of testing more hypotheses, without introducing any additional statistical assumptions. We also discuss the idea of controlling the false discovery rate for ICP.
@@ 2 A brief recap of invariant causal prediction
For the sake of simplicity, we use linear models (Peters et al.,, 2016) for illustration. Consider the following linear structural equation model:
$Y^{e}\leftarrow\beta_{1}X^{e}_{1}+\cdots\beta_{m}X^{e}_{m}+\epsilon^{e}=\beta_{S^{*}}^{T}X^{e}_{S^{*}}+\epsilon^{e},$ (1)
@@
The key idea of ICP is to test whether the conditional distribution of $Y^{e}$ is invariant across environments given predictors $X^{e}_{S}$ for some $S\subseteq[m]$. This leads to the following null hypothesis:
@@
Under the assumptions that there is no latent variable and $\mathcal{E}$ does not contain the environment where $Y$ is intervened, it is clear that $H_{0,S^{*}}(\mathcal{E})$ is true. However, there may be other sets than $S^{*}$ for which $H_{0,S}(\mathcal{E})$ is true, causing an identifiability issue. To this end, Peters et al., (2016) defined the so-called identifiable causal predictors under $\mathcal{E}$ as follows:
@@
ICP estimates the causal predictors by using the sample version of $S(\mathcal{E})$ in two steps: (i) For every $S\subseteq[m]$, test $H_{0,S}(\mathcal{E})$ at level $\alpha\in(0,1)$. (ii) Obtain the selected causal predictors by
@@
The familywise error rate guarantee of $\widehat{S}^{\text{ICP}}(\mathcal{E})$ holds at level $\alpha$ because
$P\left(\widehat{S}^{\text{ICP}}(\mathcal{E})\subseteq S^{*}\right)=P\left(\underset{S:H_{0,S}(\mathcal{E})\text{ is not rejected }}{\bigcap}S\subseteq S^{*}\right)\geq P(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected })\geq 1-\alpha.$ (5)
3 Simultaneous false discovery bounds for ICP
3.1 A multiple testing formulation of ICP
Instead of directly estimating $S(\mathcal{E})$ as in (4), another natural way is to form a multiple testing problem:
$H^{*}_{0,i}:i\in\mathcal{N}^{*},\quad i\in[m],$ (6)
where $\mathcal{N}^{*}=[m]\setminus S^{*}$. To this end, we propose the following p-value for $H^{*}_{0,i}$:
$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S},$ (7)
where $p_{S}$ is a valid p-value for testing $H_{0,S}(\mathcal{E})$ (see (2)), that is, $P_{H_{0,S}(\mathcal{E})}(p_{S}\leq c)\leq c$ for any $c\in(0,1)$. Note that it can be obtained in the first step of ICP. All p-values in this paper depend on the environment set $\mathcal{E}$, but we omit this dependence for simplicity of notation. The validity of $p^{*}_{i}$ is shown in the following proposition.
@@ 3.1 A multiple testing formulation of ICP
To connect to the original ICP approach, we need to introduce a slightly more complicated p-value. For a given $\tau\in(0,1)$, let
@@
Note that $\tilde{p}^{*}_{i}(\tau)$ is also valid because $\tilde{p}^{*}_{i}(\tau)\geq p^{*}_{i}$. Then, as the following proposition shows, the discovery set obtained by using ICP is equivalent to the one obtained by directly comparing $\tilde{p}^{*}_{i}(\alpha)$ to the significance level $\alpha$.
@@
[…]
Hence, $\tilde{S}(\mathcal{E})$ possesses the familywise error rate control guarantee (see (5)). In fact, the discovery set
@@
obtained by directly comparing $p^{*}_{i}$ to the significance level also controls the familywise error rate, as shown in the following proposition.
@@
Therefore, both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ (equivalent to $\tilde{S}(\mathcal{E})$) control the familywise error rate, and using $\widehat{S}(\mathcal{E})$ seems better because $p^{*}_{i}\leq\tilde{p}^{*}_{i}(\alpha)$. However, the scenario where $p^{*}_{i}<\tilde{p}^{*}_{i}(\alpha)$ only occurs when $p_{S}\leq\alpha$ for all $S\subseteq[m]$. In such cases, we have $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ and $\widehat{S}(\mathcal{E})=[m]$, which appears not meaningful. In practice, $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ are generally equivalent.
3.2 Simultaneous false discovery bounds for ICP
Based on the multiple testing formulation (6), we propose simultaneous false bounds by considering all intersection hypotheses. Specifically, for a given $S\subseteq[m]$, by generalizing (6) and (7), we consider null hypothesis

### v3

# On the error control of invariant causal prediction

@@ Abstract
Invariant causal prediction (ICP, Peters et al., (2016)) provides a novel way for identifying causal predictors of a response by utilizing heterogeneous data from different environments. One notable advantage of ICP is that it guarantees to make no false causal discoveries with high probability. Such a guarantee, however, can be overly conservative in some applications, resulting in few or no causal discoveries. This raises a natural question: Can we use less conservative error control guarantees for ICP so that more causal information can be extracted from data? We address this question in the paper. We focus on two commonly used and more liberal guarantees: false discovery rate control and simultaneous true discovery bound. Unexpectedly, we find that false discovery rate does not seem to be a suitable error criterion for ICP. The simultaneous true discovery bound, on the other hand, proves to be an ideal choice, enabling users to explore potential causal predictors and extract more causal information. Importantly, the additional information comes for free, in the sense that no extra assumptions are required and the discoveries from the original ICP approach are fully retained. We demonstrate the practical utility of our method through simulations and a real dataset about the educational attainment of teenagers in the US.
@@ 1 Introduction
Discovering causal predictors of the response of interest is one of the principal goals of scientific research. However, based solely on observational data, causal relationships might be unidentifiable. In such cases, heterogeneous data from different environments can enhance identifiability and thus is a valuable source for causal discovery. In the multi-environment setting, Peters et al., (2016) proposed a novel method called invariant causal prediction (ICP), which identifies causal predictors by exploiting the invariance of the conditional distribution of the response given its causal predictors across environments. Specifically, ICP generates p-values $p_{S}$ for each set $S\subseteq[m]=\{1,\dots,m\}$, where $m$ is the number of variables, by testing whether $S$ is invariant across environments (see (1)). The estimated set of causal predictors, $\widehat{S}^{\text{ICP}}$ (see (3)), is then obtained based on these p-values.
One notable advantage of ICP is that it guarantees, with high probability, that all its discoveries are true causal predictors:
$P\left(\widehat{S}^{\text{ICP}}\subseteq S^{*}\right)=P\left(|\widehat{S}^{\text{ICP}}\cap\mathcal{N}^{*}|=0\right)\geq 1-\alpha,$
@@
where $\alpha\in(0,1)$ is a nominal level, $S^{*}$ denotes the set of true causal predictors, and $\mathcal{N}^{*}=[m]\setminus S^{*}$ denotes the set of null causal predictors. This guarantee is known as familywise error rate (FWER) control.
Controlling the FWER, however, can be overly conservative in some applications, particularly when causal predictors are highly correlated with non-causal ones. In such cases, ICP may result in few or no causal discoveries. This raises the question: Is it possible to control more liberal error rates for ICP, thereby extracting more causal information?
To answer this question, we focus on two commonly used, less conservative error control guarantees: false discovery rate (FDR) control (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001) and the simultaneous true discovery bound (Genovese and Wasserman,, 2006; Goeman and Solari,, 2011). Formally, based on the p-values $p_{S}$ generated by ICP, we study the following two problems:
  • 1.
    FDR control: Can we obtain a discovery set $\widehat{S}$ satisfying the FDR control guarantee
    $\mathbb{E}\left[\frac{|\widehat{S}\cap\mathcal{N}^{*}|}{\max\{1,|\widehat{S}|\}}\right]\leq\alpha?$
    In particular, since FWER control implies FDR control (so $\widehat{S}^{\text{ICP}}$ already guarantees FDR control), the goal here is not merely to obtain a set $\widehat{S}$ with FDR control, but also to have $\widehat{S}$ contain more discoveries than $\widehat{S}^{\text{ICP}}$.
  • 2.
@@
    Simultaneous true discovery bound: Can we obtain a function $t_{\alpha}:2^{[m]}\rightarrow[0,1]$, where $2^{[m]}$ denotes the power set of $[m]$, such that
    $P(|R\cap S^{*}|\geq t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha?$
    This bound enables users to freely check any set $R$ they consider interesting, which may be formed based on applying some machine learning methods or their expertise, with $t_{\alpha}(R)$ providing a high probability lower bound on the number of true discoveries in $R$.
To address these two problems, our main idea is to consider a multiple testing problem (see (4)) that directly tests whether causal predictors are null, and propose p-values based on $p_{S}$ generated by ICP. We can then apply false discovery rate control procedures or construct simultaneous true discovery bounds using these p-values. To this end, we propose valid p-values $p_{i}^{*}$ (see (5)). Furthermore, we prove that the original ICP method is equivalent to directly comparing certain p-values (see (6)), which are slightly more conservative than $p_{i}^{*}$, to the significance level.
For the first problem, our answer is negative. We find that directly comparing $p_{i}^{*}$ to the nominal level already controls the FWER, which is somewhat uncommon because multiple testing procedures, such as Bonferroni’s correction, are typically required to control the FWER. Given this, there is no reason to further applying FDR control procedures to $p_{i}^{*}$, and the only hope for obtaining more causal discoveries through FDR control would be if $p^{*}_{i}$ were conservative. If that were the case, we may propose better p-values and obtain more causal discoveries by applying FDR control methods to them. As a result, we investigate whether $p_{i}^{*}$ is conservative in a special setting where the optimal $p_{S}$ can be obtained by ICP. We show that $p_{i}^{*}$ is not conservative in this setting, indicating that obtaining smaller valid p-values is unlikely. Consequently, we believe that it is unlikely to obtain more causal discoveries by controlling the FDR for ICP.
For the second problem, our answer is positive. The multiple testing framework provides a convenient way to apply the closed testing procedure from Goeman and Solari, (2011) to obtain simultaneous true discovery bounds. Furthermore, we show that due to the special construction of $p_{i}^{*}$, the closed testing adjustment step is unnecessary, which significantly reduces computation as we do not need to implement all steps of the standard closed testing procedure. We also present a simultaneous bound based on the original ICP formulation, and we prove its equivalence to the one obtained by closed testing. This connection is useful as it provides a better understanding of the statistical properties, such as admissibility, of the simultaneous true discovery bound. These properties are less evident in the original ICP formulation but have been well-studied in the context of closed testing (Goeman et al.,, 2021).
@@
The simultaneous true discovery bound preserves the same causal information obtained by the original ICP approach, and it can extract additional causal insights. Hence, it is always beneficial to use it in practice. As a quick illustration, for a simulated dataset with nine predictors shown in Section 5.1, no information is obtained by the original ICP approach as it returns zero discoveries. However, more causal information can be extracted by using the simultaneous true discovery bounds. For example, we can know that with a probability larger than $0.95$, there is at least one causal predictor in $\{X_{3},X_{7}\}$, at least two causal predictors in $\{X_{3},X_{5},X_{7},X_{8}\}$, and at least four causal predictors in all nine predictors. More simulations and a real data application can be found in Section 5.
We note that the simultaneous true discovery bound $t_{\alpha}(R)$ is equivalent to the simultaneous false discovery bound, in the sense that based on $t_{\alpha}(R)$, we can obtain simultaneous false discovery bound $|R|-t_{\alpha}(R)$ satisfying
$P(|R\cap\mathcal{N}^{*}|\leq|R|-t_{\alpha}(R)\text{ for any }R\subseteq[m])\geq 1-\alpha.$
We focus on the simultaneous true discovery bound throughout this paper because it offers a more convenient interpretation for applications. Many other error rates, such as k-FWER (Lehmann and Romano,, 2005; Sarkar,, 2007) and false discovery exceedance (Dudoit et al.,, 2004; Korn et al.,, 2004; Romano and Shaikh,, 2006), can be seen as special cases of the simultaneous true discovery bound, see Goeman et al., (2021) for further discussions.
At last, Heinze-Deml et al., (2018) also investigated the question of mitigating the conservativeness of FWER control for ICP, and they proposed the so-called defining sets to tackle this problem. Defining sets are the smallest sets guaranteed to contain at least one true causal predictor, making them particularly useful when ICP returns an empty set. As we will see later, the true discovery bound encompasses defining sets as special cases.
@@ 2 A brief recap of invariant causal prediction
For the sake of simplicity, we use linear models (Peters et al.,, 2016) for illustration. Our results can also be applied to non-linear cases (Heinze-Deml et al.,, 2018) directly. Consider the following linear structural equation model:
$Y^{e}\leftarrow\beta_{1}X^{e}_{1}+\cdots\beta_{m}X^{e}_{m}+\epsilon^{e}=\beta_{S^{*}}^{T}X^{e}_{S^{*}}+\epsilon^{e},$
@@
The key idea of ICP is to test whether the conditional distribution of $Y^{e}$ is invariant across environments given predictors $X^{e}_{S}$ for each $S\subseteq[m]$. This leads to the following null hypothesis:
@@
Under the assumptions that there is no latent variables and $\mathcal{E}$ does not contain the environment where $Y$ is intervened, it is clear that $H_{0,S^{*}}(\mathcal{E})$ is true. However, there may be other set $S$ than $S^{*}$ for which $H_{0,S}(\mathcal{E})$ is true, causing an identifiability issue. To this end, Peters et al., (2016) defined the so-called identifiable causal predictors under $\mathcal{E}$ as follows:
@@
ICP estimates the causal predictors by using the sample version of $S(\mathcal{E})$ in two steps:
  • (i)
    For each $S\subseteq[m]$, test $H_{0,S}(\mathcal{E})$ at level $\alpha\in(0,1)$ to obtain p-value $p_{S}$.
  • (ii)
    Obtain the estiamted causal predictors by
@@
The FWER guarantee of $\widehat{S}^{\text{ICP}}(\mathcal{E})$ holds at level $\alpha$ because
$P\left(\widehat{S}^{\text{ICP}}(\mathcal{E})\subseteq S^{*}\right)=P\left(\underset{S:H_{0,S}(\mathcal{E})\text{ is not rejected }}{\bigcap}S\subseteq S^{*}\right)\geq P(H_{0,S^{*}}(\mathcal{E})\text{ is not rejected })\geq 1-\alpha.$
3 A multiple testing formulation of ICP
The original formulation of ICP described in Section 2, while elegant and concise, may not be the most suitable for addressing the two problems described in Section 1. In this section, we provide a multiple testing perspective on the original ICP approach, which offers additional insights for addressing the considered problems.
Specifically, instead of directly estimating $S(\mathcal{E})$ as in (3), another natural approach is to consider the following multiple testing problem:
$H^{*}_{0,i}:i\in\mathcal{N}^{*},\quad i\in[m].$ (4)
Recall that $\mathcal{N}^{*}=[m]\setminus S^{*}$ is the set of null causal predictors. The idea is to obtain valid p-values for these hypotheses, then apply false discovery rate control procedures, such as the Benjamini-Yekutieli procedure (Benjamini and Yekutieli,, 2001), or construct simultaneous true discovery bounds using these p-values. These approaches may lead to more discoveries compared to $\widehat{S}^{\text{ICP}}(\mathcal{E})$, because they are less conservative than controlling the FWER.
To this end, based on the p-values $p_{S}$ generated by the first step of ICP, we propose the following p-value for $H^{*}_{0,i}$:
$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}.$ (5)
All p-values in this paper depend on the environment set $\mathcal{E}$, but we surpress this dependence for simplicity in notation. We take the p-values $p_{S}$ as given from the ICP approach and do not elaborate on their calculation. For more details, please refer to Peters et al., (2016).
As shown in the following proposition, $p^{*}_{i}$ is a valid p-value.
@@ 3 A multiple testing formulation of ICP
One may wonder how the original ICP method connects to this multiple testing formulation, and we now establish this connection. Since the original ICP method controls the FWER, one might guess that applying some FWER control procedure, such as the Bonferroni correction or Holm’s procedure (Holm,, 1979), would yield a set equivalent to $\widehat{S}^{\text{ICP}}(\mathcal{E})$. However, this is not the case. Perhaps surprisingly, $\widehat{S}^{\text{ICP}}(\mathcal{E})$ is almost equivalent to the set obtained by directly comparing $p^{*}_{i}$ to the nominal level, without applying any multiple testing corrections.
We need to introduce a slightly more complicated p-value than $p^{*}_{i}$ in order to establish the connection. For a given $\tau\in(0,1)$, let
@@
Note that $\tilde{p}^{*}_{i}(\tau)$ is also a valid p-value because $\tilde{p}^{*}_{i}(\tau)\geq p^{*}_{i}$. Then, as the following proposition shows, the discovery set $\widehat{S}^{\text{ICP}}(\mathcal{E})$ obtained by ICP is equivalent to the set obtained by directly comparing $\tilde{p}^{*}_{i}(\alpha)$ to the nominal level $\alpha$.
@@
For $\alpha\in(0,1)$, let $\tilde{S}(\mathcal{E})=\{i:\tilde{p}^{*}_{i}(\alpha)\leq\alpha\}$. Then $\tilde{S}(\mathcal{E})=\widehat{S}^{\text{ICP}}(\mathcal{E})$.
As a result, $\tilde{S}(\mathcal{E})$ possesses the FWER control guarantee. In fact, the discovery set
@@
obtained by directly comparing $p^{*}_{i}$ to the nominal level $\alpha$ also controls the FWER, as shown in the following proposition.
@@
Note that $\widehat{S}^{\text{ICP}}(\mathcal{E})=\tilde{S}(\mathcal{E})\subseteq\widehat{S}(\mathcal{E})$ as $p^{*}_{i}\leq\tilde{p}^{*}_{i}(\alpha)$. Given that both $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ control the FWER, using $\widehat{S}(\mathcal{E})$ seems to improve upon the original ICP approach. However, $p^{*}_{i}<\tilde{p}^{*}_{i}(\alpha)$ only occurs when $p_{S}\leq\alpha$ for all $S\subseteq[m]$. In such cases, $\widehat{S}^{\text{ICP}}(\mathcal{E})=\emptyset$ and $\widehat{S}(\mathcal{E})=[m]$, which seems useless because even $p_{S^{*}}\leq\alpha$, leading to a false rejection of $H_{0,S^{*}}(\mathcal{E})$. In practice, $\widehat{S}(\mathcal{E})$ and $\widehat{S}^{\text{ICP}}(\mathcal{E})$ are generally equivalent.
4 Can we control less conservative errors for ICP?
4.1 FDR control does not seem well-suited for ICP
Proposition 3.3 shows that directly comparing $p^{*}_{1},\dots,p^{*}_{m}$ to the nominal level automatically ensures FWER control, as if there were no multiple testing issue. This is somewhat unexpected, as one would typically need to apply multiple testing adjustments (e.g., Bonferroni correction) to control the FWER. Consequently, there is no need to further apply FDR control procedures to $p^{*}_{1},\dots,p^{*}_{m}$.
Given this, the only remaining hope for obtaining more causal discoveries through FDR control is that the construction (5) of $p^{*}_{i}$ based on $p_{S}$ is conservative. Then, if a more powerful method for obtaining smaller p-values for $H^{*}_{0,i}$ based on $p_{S}$ can be proposed, we might return to the “normal” situation where directly comparing these p-values to the nominal level no longer controls the FWER. If this could be achieved, applying FDR control procedures may lead to more causal discoveries. Unfortunately, as we will show below, $p^{*}_{i}$ defined by (5) does not appear to be conservative.
We consider a special case where the optimal $p_{S}$ can be obtained for testing $H_{0,S}(\mathcal{E})$. That is, we have $p_{S}\sim U(0,1)$ for true null hypothesis and $p_{S}=0$ for false null hypothesis. This represents the most desirable scenario, where powerful tests for $H_{0,S}(\mathcal{E})$ are available. It is also called the Dirac-uniform configuration (Finner et al.,, 2009). Note that obtaining these $p_{S}$ values is a separate task from constructing a valid p-value for the hypothesis (4) based on them.
Then, for $i\in S(\mathcal{E})$ (see (2)), which is an identifiable causal predictor based on ICP, the construction (5) yields
$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}=0,$
because $H_{0,S}(\mathcal{E})$ is false for any set $S$ that does not contain $i$, implying that $p_{S}=0$. In other words, for any identifiable causal predictor, the corresponding p-value will be 00, which is the best possible outcome. Therefore, (5) does not seem to be a conservative construction for $p^{*}_{i}$.
Furthermore, if we are willing to additionally assume the following weak Assumption 4.1, then in the case where the environment set is sufficiently rich, we can show that $p^{*}_{i}$ is admissible, as demonstrated below.
Assumption 4.1 (Common null events).
For every measurable event $E$ and every possible data distribution $P_{1}$ and $P_{2}$, if $P_{1}(E)=0$, then $P_{2}(E)=0$.
This assumption means that the collection of null events is common across all considered possible distributions. It is satisfied by most models, provided that some deterministic corner cases are excluded. For more discussion, see the Appendix A of Goeman et al., (2021).
In addition, assume that we are in a scenario where the only true null hypothesis is $H_{0,S^{*}}(\mathcal{E})$. This is the desirable setting in which all causal predictors are identifiable, and it can occur, for example, when the environment set is sufficiently rich. In this scenario, for any $i$ such that $H^{*}_{0,i}$ is true, we have
$p^{*}_{i}=\max_{S\subseteq[m]\setminus\{i\}}p_{S}=p_{S^{*}}.$
In this special case, all null p-values are equal to $p_{S^{*}}$, providing a scenario where it is clear that directly comparing the p-values to the nominal level controls the FWER. Then, we have
$P_{H^{*}_{0,i}}(p^{*}_{i}\leq\alpha)=P_{H^{*}_{0,i}}(p_{S^{*}}\leq\alpha)=\alpha.$
By the following Lemma 4.1, $p^{*}_{i}$ is admissible. Note that both obtaining optimal $p_{S}$ and having a sufficiently rich environment set $\mathcal{E}$ are independent of the underlying truth regarding which variables are true causal predictors. The admissibility $p^{*}_{i}$ suggests that its construction does not seem to be conservative.
Lemma 4.1 (Lemma $7$ in the Appendix A of Goeman et al., (2021)).
Under Assumption 4.1, a p-value $p$ of hypothesis $H$ is admissible if there exist a distribution $P$ under $H$ such that $P_{H}(p\leq\alpha)=\alpha$.
In general settings where $p_{S}$ is not optimal, we can not fully exclude the possibility of constructing a more powerful (i.e., smaller) $p^{*}_{i}$. However, based on the discussions above, we believe it is unlikely that $p^{*}_{i}$ can be significantly improved to yield more discoveries after applying FDR control procedures to the new p-values. Therefore, our response to the first question raised in the introduction is negative: FDR control does not seem well-suited for ICP, and the idea of obtaining more discoveries by controlling the FDR in the context of ICP appears unpromising.
4.2 Simultaneous true discovery bound for ICP
Now, we address the second question about simultaneous true discovery bound. Fortunately, this time, the answer is positive: we can obtain a simultaneous true discovery bound and use it to extract more causal information. Below, we present how to obtain this bound.
Based on the multiple testing formulation (4), a straightforward way to obtain simultaneous true discovery bound is to apply the closed testing procedure (Marcus et al.,, 1976; Goeman and Solari,, 2011) (see Appendix A for a brief recap). This requires testing the intersection hypothesis